Grade 12th passMechanics

The resultant of two vectors at an angle 150 degree is 10 units and is perpendicular to one vector. The magnitude of the smaller vector is

Haider bohri
8 Years agoGrade 12th pass

3 Answers

Harsh jain
8 Years ago
the ans to this is very easy and straightforward...let two vectors v1 and v2..v2 sin60 = v1...and also v2 cos60 = 10 ......So v2 = 20 ......and v1 = 10 root 3
Jahnvi Pandey
8 Years ago
In x direction
Bcos 60°=10
Since cos 60°=1/2
B=10/1/2
=20
In y direction
A=B sin 60°
Since sin 60°=√3/2
A=20√3/2= 10√3
Since 10√3 is smaller than 20.
Kushagra Madhukar
5 Years ago
Dear student,
Please find the solution to your problem below.
 
Let the two vectors be A(along the x -axis), B(150o from x-axis in anticlockwise sense, i.e. 60o from y-axis)
Let the resultant be perpendicular to A, thus it will be along y-axis with a magnitude, R =10
 
Now, In y direction
Bcos 60° = R
Since cos 60°=1/2
B=10/0.5
=20
 
In x direction
A – B sin 60° = 0
A = B sin 60°
Since sin 60°=√3/2
A=20√3/2= 10√3
 
Hence the smaller vector has a magnitude of 10√3
 
Hope it helps.
Thanks and regards,
Kushagra